OpenStudy (watchmath):

Find the solution of the following system of equations \[ \begin{align*} \frac{1}{x}+\frac{1}{y+z}&=\frac{1}{2}\\ \frac{1}{y}+\frac{1}{x+z}&=\frac{1}{3}\\ \frac{1}{z}+\frac{1}{x+y}&=\frac{1}{4} \end{align*} \]

OpenStudy (anonymous):

Hm

OpenStudy (anonymous):

well i spent 10 minutes typing my solution to this and my browser crashed. I'll give my result and the basic idea: x=23/10 y=23/6 z=23/2 we can write the system as: 2(x+y+z)=x(y+z) 3(x+y+z)=y(x+z) 4(x+y+z)=z(x+y) isolating x+y+z on the LHS of each we get the equality: (xy+xz)/2=(xy+zy)/3=(xz+zy)/4 from here we find that: y=z/3 x=z/5 plugging these into the first given eqution we find z=23/2 and then x and y are found easily.